# Multi-scale Self-similar Finite-time Blowups of the Constantin–Lax–Majda Model for the 3D Euler Equations

### v2

@@ 3.1. Example I: Exact one-scale self-similar blowup
We first investigate the exact self-similar solution by choosing
@@
3.2. Example II: One-scale asymptotically self-similar blowup
@@
We use (3.3) to find that
@@
$Z(t)=\begin{cases}\displaystyle\pm\frac{1}{2}\sqrt{1-\frac{25}{16}t^{2}}+\mathrm{i}\left(\frac{5}{8}t-1\right),&t\in[0,\frac{4}{5}]\\ \\ \displaystyle\mathrm{i}\left(\pm\frac{1}{2}\sqrt{\frac{25}{16}t^{2}-1}+\frac{5}{8}t-1\right),&t\in(\frac{4}{5},1].\end{cases}$
@@
travels in curved lines from their starting points to the joining point $-\mathrm{i}/2$ where they merge into one point at $t=4/5$. Then, in the second phase $t\in[4/5,1]$, the two zeros,
@@ 3.3. Example III: Two-scale blowup for $n=1$
We use (3.3) to find that
@@
The trajectory of $Z_{1}(t)$ is not a straight line, which explains why in this case the solution $\omega(x,t)$ is not exactly self-similar. Moreover, we can see that $|Y_{1}(t)|\sim(1-t)$ converges to 00 much faster than $|X_{1}(t)|\sim(1-t)^{1/2}$. This verifies the two-scale blowup phenomenon described in Theorem 2.2.
@@
which verifies the asymptotic formula of $\omega(x,t)$ in Theorem 2.2 with $a=1$, $b=1/2$, $c=1/4$.
@@ 3.6. Example VI: Non-blowup traveling wave
As we have argued in the previous section (see (2.9) with $c=1$, $r=-1$), the solution to (1.1) with this initial data $\omega_{0}$ is a traveling wave
@@ 3.7. General analysis
As we have mentioned, analyzing the evolution of all poles of $\eta(z,t)$ for general initial data $\eta_{0}$ can be very difficult. Nevertheless, we can still perform a relatively general analysis on the local behavior of poles of $\eta(z,t)$ as they approach the real line $\mathbb{R}$.
Let us write
@@
for some $a,c>0$, $b\in\mathbb{R}$ and some integer $n\geq 0$. We also assume that $b>0$ when $k\geq 1$. Note that $n=0$ corresponds to the one-scale case in Theorem 2.1, $n=1$ corresponds to the basic two-scale case in Theorem 2.2, and $n\geq 1$ corresponds to the general two-scale case in Theorem 2.3.

### v3

@@ 3.3. Example I: Exact one-scale self-similar blowup
Let us first choose
@@
3.4. Example II: Asymptotic one-scale self-similar blowup
@@
We use (3.3) to find that the zeros of $\zeta(z,t)$ satisfy
@@
$Z(t)=\begin{cases}\displaystyle\pm\frac{1}{2}\sqrt{1-\frac{25}{16}t^{2}}+\mathrm{i}\left(\frac{5}{8}t-1\right),&t\in[0,\frac{4}{5}],\\ \\ \displaystyle\mathrm{i}\left(\pm\frac{1}{2}\sqrt{\frac{25}{16}t^{2}-1}+\frac{5}{8}t-1\right),&t\in(\frac{4}{5},1].\end{cases}$
@@
travel in curved lines from their starting points to the joining point $-\mathrm{i}/2$ where they merge into one point at $t=4/5$. Then, in the second phase $t\in[4/5,1]$, the two zeros,
@@ 3.5. Example III: Two-scale blowup for $n=1$
We use (3.3) to find that the zeros of $\zeta(z,t)$ satisfy
@@
The trajectory of $Z_{1}(t)$ is not a straight line, which explains why in this case the solution $\omega(x,t)$ is not exactly self-similar. Moreover, we can see that $|Y_{1}(t)|\sim(1-t)$ converges to 00 much faster than $|X_{1}(t)|\sim(1-t)^{1/2}$. This agrees with the two-scale blowup phenomenon described in Theorem 2.2.
@@
which agrees with the asymptotic formula of $\omega(x,t)$ in Theorem 2.2 with $a=1$, $b=1/2$, $c=1/4$.
@@ 3.8. Example VI: Non-blowup traveling wave
As we have argued in the previous section (see (2.10) with $c=1$, $r=-1$), the solution to (1.1) with initial data $\omega_{0}$ is a traveling wave
@@ 3.9. General analysis
As we have mentioned, analyzing the evolution of all poles of $\eta(z,t)$ for general initial data $\eta_{0}$ can be very difficult. Nevertheless, we can still perform a relatively general analysis on the local behavior of poles of $\eta(z,t)$ as they approach the real line $\mathbb{R}$. Let us write
@@
for some $a,c>0$, $b\in\mathbb{R}$ and some integer $n\geq 0$. We also assume that $b>0$ when $n\geq 1$. Note that $n=0$ corresponds to the one-scale case in Theorem 2.1, $n=1$ corresponds to the basic two-scale case in Theorem 2.2, and $n\geq 1$ corresponds to the general two-scale case in Theorem 2.3.